OpenStudy (anonymous):

find (fog) (2) for f(x)= -7x+3 and g(x) =x^2+4

OpenStudy (anonymous):

f(x)=??? the entry is wrong. @Arizona1

OpenStudy (anonymous):

f o g(x) Is a composition of functions I think you mistyped your f(x) Because f(x) has no x dependance, f will always be -4, and so f o g will always be -4

OpenStudy (anonymous):

my keys still shows no to both answers, please help :(

OpenStudy (anonymous):

sorry I'm an idiot wrote it wrong sorry

OpenStudy (anonymous):

(fog) (2) for f(x)= -7x+3 and g(x) =x^2+4 F comp g of x = -7[x^2+4]+3 = -7x^2-28+3=-7x^2-25

OpenStudy (anonymous):

@Arizona1

OpenStudy (anonymous):

(fog) (x)=f(g(x))=f(x^2+4)=-7(x^2+4)+3 put x=2 and we have -7(4+4)+3=-56+3=-53